OpenStudy (anonymous):

Help Algebra is not my best I need to solve this L =P-2W __________ 2 ?

jimthompson5910 (jim_thompson5910):

Solve for what?

OpenStudy (anonymous):

Part A: Solve for P. After solving this formula for P, explain the use of the formula. Part B: Next, create your own values for W and L. Compute your value for P with your own data.

jimthompson5910 (jim_thompson5910):

\[\Large L = \frac{P - 2W}{2}\] \[\Large 2L = P - 2W\] \[\Large 2L + 2W = P\] \[\Large P = 2L + 2W\] For part B, plug in any two numbers you want for L and W. Then evaluate the right side to find P.

OpenStudy (anonymous):

wow that took you a couple of seconds I have been trying how about a rate and distance problem?

jimthompson5910 (jim_thompson5910):

sure go ahead and ask

OpenStudy (anonymous):

The number of miles remaining for a jet traveling from San Francisco to New York can be approximated with the linear equation Miles remaining = 2650 - 475t where t is the number of hours since leaving San Francisco. Part A: In how many hours will the jet be 1,000 miles from New York? Part B: How long will the whole trip take?

jimthompson5910 (jim_thompson5910):

m = miles remaining m = 2650 - 475t 1000 = 2650 - 475t 1000-2650 = -475t -1650 = -475t -1650/(-475) = t 3.47368 = t t = 3.47368 So it will be roughly 3.47368 hours when the jet is 1000 miles from new york For part b), plug in m = 0 and solve for t.

OpenStudy (anonymous):

A runner's heart rate in beats per minute is given by the formula Heart rate = 65 + 53 / (4t + 1) where t is the number of minutes after the start of a cool-down period. Part A: Find the runner's heart rate in 10 minutes. Part B: When is the runner's heart rate equal to 68 beats per minute? Does this use the same formula as rate and distance?

jimthompson5910 (jim_thompson5910):

No, this is an entirely different problem. But you follow the same basic steps to solve the equation. part A) plug in t = 10 and evaluate/simplify part B) set the entire right side equal to 68 and solve for t

OpenStudy (anonymous):

Gosh thank you I have been struggling as I said its not my best but you really made my day, in just a few minutes I am impressed...

jimthompson5910 (jim_thompson5910):

I'm glad I could help you out

jimthompson5910 (jim_thompson5910):

Part B: Next, create your own values for W and L. Compute your value for P with your own data. This part?

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

Alright, which values did you pick for L and W?

OpenStudy (anonymous):

5 and 10

jimthompson5910 (jim_thompson5910):

Alright, so plug them into the equation P = 2L + 2W what do you get?

OpenStudy (anonymous):

15?

jimthompson5910 (jim_thompson5910):

P = 2L + 2W P = 2(10) + 2(5) ....simply replace L with 10 and W with 5

jimthompson5910 (jim_thompson5910):

then what?

OpenStudy (anonymous):

subtract 2 from each

jimthompson5910 (jim_thompson5910):

2(10) means 2 times 10

jimthompson5910 (jim_thompson5910):

so we multiply each 10 and 5 by 2

OpenStudy (anonymous):

then since it is + subtract?

jimthompson5910 (jim_thompson5910):

no, simply add since you're not isolating any terms

OpenStudy (anonymous):

so p=30

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

do you have a tutoring sight?

jimthompson5910 (jim_thompson5910):

or you can email me if you like (it's on the site) Alternatively, you can use msn or yahoo messenger to get in contact with me (my screen name is the same as my name here)

jimthompson5910 (jim_thompson5910):

Did you have any other questions?

OpenStudy (anonymous):

I will be back I am on the West Coast you? I am registered at algebra.com as of a few minutes ago I am failing algebra and its part of my associates, Just need to understand Algebra, Its hard to teach a OLD dog new tricks..LOL

jimthompson5910 (jim_thompson5910):

lol glad you signed up there as I tutor over there as well (small world lol), feel free to email me or add me on msn or yahoo messenger oh and yes, I'm in CA, so I'm also on the west coast

OpenStudy (anonymous):

Great thanks

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Hey Jim_thompson5910 are you still on here

OpenStudy (anonymous):

. A cell phone company has offered you two different rates. The first rate is $35 per month and 5 cents per phone call (time does not matter). The second offer is $55 per month and 2 cents per call. Part A: Set up and solve the inequality to model this problem and find where it is less expensive to go with the second plan. The cell phone company also has two texting plans. The first is unlimited texts for $12. The second plan is 8 cents per text. Part B: Set up and solve an inequality to show when it less expensive to go with the second texting plan.